A particle is thrown up with the velocity of 40m/s. The maximum height reached by the particle is?

Grade:9

2 Answers

salmankhan
6 years ago
formula of maximum range is 
Rmax =U/g
substituting the value 
=40*40/10
=1600/10
=160 m
 
…...............................................................................................................
Janhavi
6 years ago
U^2sin^2theata/2g=height attainedTheata=90 i.e=1 and g=10(40)^2×1/2×10= max height attained 1600/20=80So max height attained by particle is 80 m.